# Steering Language Model to Stable Speech Emotion Recognition via Contextual Perception and Chain of Thought

### v1

@@ Abstract
[…]
booktabs
Steering Language Model to Stable Speech Emotion Recognition via Contextual Perception and Chain of Thought
[…]
Figure 1: Overview of C2SER. C2SER overcomes the issue of hallucinations through contextual perception and chain of thought.
1 Introduction
@@
Audio is a multifaceted medium for communication, conveying speech prosody, vocal tone, and paralinguistic cues through its acoustic features. Large-scale audio-language models (ALMs) Wang et al. (2024) have demonstrated substantial progress in understanding diverse forms of audio signals, which is crucial for advancing Artificial General Intelligence (AGI). With increasing data availability, computational power, and model size, significant strides have been made in speech signal comprehension, analysis, and reasoning, leading to more natural and human-like text responses.
Speech emotions are crucial in communication, influencing how individuals interact and respond through variations in tone, rhythm, and intensity. However, ALMs still face challenges in speech emotion recognition (SER) Bellver-Soler et al. (2024); Akman et al. (2025), which aims to accurately detect emotional states from spoken language. Speech emotions are inherently complex, encompassing diverse affective states, and their precise recognition is essential for enhancing communication effectiveness across diverse domains. Despite advancements in general audio understanding, ALMs often struggle with hallucinations in the SER task due to the intricate context of speech. These hallucinations lead to misclassifications, where the model may incorrectly interpret the emotional tone or generate irrelevant responses, undermining the model’s reliability in real-world applications.
@@
In this work, we address the hallucination problem in ALM-based SER by incorporating detailed speech information and expanding the model’s reasoning length and depth. As illustrated in Figure 1, we introduce C2SER, a novel ALM designed to improve both the stability and accuracy of SER. C2SER integrates two critical components: contextual perception and a chain of thought (CoT), leveraging both speech content and speaking styles (e.g., speaking rate, pitch, energy) to facilitate emotion recognition. The encoder of Whisper Radford et al. (2023), trained for automatic speech recognition (ASR), speech translation, and language identification, is employed for semantic perception. For acoustic perception, we introduce Emotion2Vec-S, a refined extension of Emotion2Vec Ma et al. (2024b), designed to enhance the extraction of emotion-related information from audio. Emotion2Vec-S incorporates semi-supervised contrastive loss at the category level, improving emotional discrimination by combining self-supervised and semi-supervised learning.
@@
Recognizing that speech emotions are influenced by both speech content and speaking styles, such as aggressive speech characterized by a loud volume potentially indicating anger, C2SER employs a CoT Guo et al. (2025) training approach to incentivize reasoning capability. This approach decomposes the SER task into sequential steps: first perceiving speech content and speaking style, followed by emotion inference, with the assistance of prior context. This structured method imitates human thinking and reduces the possibility of hallucinations. To further enhance stability and prevent error propagation, especially in longer thought chains, C2SER introduces self-distillation, transferring knowledge from explicit to implicit CoT. This process helps minimize error accumulation, improving the model’s overall performance.
@@
We validate the effectiveness of C2SER through extensive experiments using multiple speech corpora and compare C2SER with state-of-the-art models, including SECap Xu et al. (2024) and Qwen2-Audio Chu et al. (2024). To better simulate real-world context, we introduce a new SER test set, Emo-Emilia. Emo-Emilia is created through an automated labeling approach on the in-the-wild Emilia corpus and manually verified to ensure quality and diversity across various scenarios. Our experimental results demonstrate that C2SER significantly outperforms existing models on public test sets and Emo-Emilia in terms of weighted accuracy (WA), unweighted accuracy (UA), and Macro F1 score, while notably reducing hallucination-related errors. These findings highlight the potential of C2SER to provide stable and reliable emotion recognition in diverse contexts.
@@
  • We release the code, checkpoints, and test sets to promote further research in the field of SER.
2 Related Work
2.1 Audio Language Model
@@
Large language models (LLMs) have made significant strides in natural language processing (NLP), showcasing remarkable capabilities across a variety of tasks Zhao et al. (2023). As audio is a critical medium of communication in human interactions and human-computer engagement, recent research has extended LLMs to integrate the audio modality, leading to the development of audio-language models (ALMs) Wang et al. (2024). ALMs tackle tasks such as audio event detection, audio captioning, and speech recognition, serving as a cornerstone for comprehensive audio understanding Chu et al. (2023).
@@
With the rapid advancements in both LLMs and the audio domain, ALMs have gained significant attention for their powerful general audio comprehension abilities. A typical ALM architecture consists of three core components: an audio encoder for modality-specific feature extraction, an LLM for text generation, and a projection layer to bridge the gap between the audio and text modalities. In addition to these foundational components, several studies have focused on refining ALM performance through innovative model architectures. For example, SALMONN Tang et al. (2024) utilizes dual encoders to separately process speech and non-speech audio signals, effectively mitigating potential conflicts between different types of audio input. Other approaches have explored training strategies to enhance ALMs’ capabilities, with Qwen2-Audio Chu et al. (2024) being a notable example. This model employs a comprehensive training pipeline that includes pretraining, supervised fine-tuning (SFT), and reinforcement learning from human feedback (RLHF).
@@
While significant progress has been made in improving the generalization and intelligence of ALMs, their performance in speech emotion recognition (SER) remains unsatisfactory, primarily due to hallucinations. SER is particularly challenging because speech emotions are inherently complex and context-dependent Yang et al. (2024), making it difficult for ALMs to interpret emotional states accurately. In many cases, the model may be misled by the content of the speech, leading to incorrect classifications or irrelevant responses.
2.2 Speech Emotion Recognition
@@
Speech emotion is a key form of paralinguistic information that effectively conveys the speaker’s intent. Speech Emotion Recognition (SER) aims to automatically identify the emotional state of a speaker from their speech. The typical SER pipeline consists of three stages: speech preprocessing, feature extraction, and emotion classification Khare et al. (2024). Early studies relied on manually engineered feature sets, such as MFCC, and simple neural network architectures like CNN and RNN, achieving basic performance on laboratory datasets (e.g., CREMA-D Cao et al. (2014), IEMOCAP Busso et al. (2008)).
@@
To address the challenge of recognizing diverse emotional expressions in real-world environments, recent research has shifted towards self-supervised learning (SSL) models Naini et al. (2024); Zhu and Sato (2023), known for their powerful generalization capabilities. SSL models are trained on large-scale unlabeled speech data in an unsupervised manner, allowing them to extract rich, generalizable representations directly from raw speech waveforms. Popular SSL models, such as HuBERT Hsu et al. (2021) and wav2vec 2.0 Baevski et al. (2020), have demonstrated significant effectiveness in extracting emotional features, serving as robust encoders for SER tasks. Additionally, researchers have explored emotion-specific SSL models designed to capture emotion-relevant features. A popular approach involves fine-tuning SSL models on emotionally labeled data for specific emotional tasks. A prominent example is Emotion2Vec Ma et al. (2024b), which is pre-trained on emotional data through self-supervised online distillation. Emotion2Vec uses both utterance-level and frame-level loss as supervision, demonstrating remarkable improvements in emotion recognition across different languages.
@@
3 Method
Figure 2: Detailed structure of C2SER. C2SER leverages speech content and speaking styles through chain-of-thought to assist speech emotion recognition.
3.1 Framework Overview
@@
More specifically, the contextual perception module comprises the following elements: a Whisper Radford et al. (2023) encoder for semantic perception, Emotion2Vec-S for acoustic perception, and a connection model designed to align the feature dimensions with those required by the text LLM. Formally, given a speech waveform $X$, the Whisper encoder extract semantic representations $S={s_{1},s_{2},...,s_{N}}$ and the Emotion2Vec-S extracts acoustic representations $A={a_{1},a_{2},...,a_{M}}$ from $X$. Let $Y={y_{1},y_{2},...,y_{T}}$ be the text descriptions and $P={p_{1},p_{2},...,p_{L}}$ be the text prompts. The text LLM, parameterized by $\theta$, takes $S$ and $A$ as input and predicts $Y$ in an autoregressive manner. The overall process can be formulated as a conditional probability:
@@ 3.2 Contextual Perception
Emotion2Vec-S is built upon the universal speech emotion representation model, Emotion2Vec. Emotion2Vec follows the architecture of data2vec Baevski et al. (2022), consisting of several standard Transformer blocks, and is pre-trained on open-source, unlabeled emotion data using self-supervised online distillation. It combines an utterance-level loss to capture global emotional features with a frame-level loss to capture contextual emotion nuances. Emotion2Vec has demonstrated superior performance compared to previous state-of-the-art pre-trained universal models and emotion specialist models, which often require only the training of linear layers for the SER task on emotional speech corpora. However, because Emotion2Vec focuses solely on the emotional expressions within individual utterances, it potentially struggles to distinguish similar utterances that express different emotions, such as fear and sadness.
With the above observation, Emotion2Vec-S introduces a coarse-level supervision to Emotion2Vec. Vanilla Emotion2Vec expands data2vec2.0 Baevski et al. (2023) with a fixed number of utterance tokens and is trained with $\mathcal{L}_{Utt}$ to learn the global emotion and $\mathcal{L}_{Frm}$ to learn the context emotion. Inspired by CLIP Radford et al. (2021), Emotion2Vec-S extends Emotion2Vec with a category-level contrastive loss $\mathcal{L}_{Cate}$. Specifically, let $G$ be the global embedding of Emotion2Vec after average pooling. Emotion2Vec-S applies a contrastive loss on $G$ by treating embeddings from utterances of the same emotion category as positive pairs and those from different categories as negative pairs. The model calculates cosine similarities between these embeddings, maximizing the similarity of positive pairs while minimizing that of negative pairs. The overall loss function for Emotion2Vec-S is formulated as follows:
@@
3.3 Explicit Chain-of-Thought
Explicit CoT reasoning enhances the ability of LLMs to handle specific tasks by detailing intermediate steps, thereby guiding the model through its reasoning process Yue et al. (2024); Yu et al. (2024). In C2SER, explicit CoT is employed to sequentially address the SER task. After the contextual perception module extracts detailed information regarding speech content and speaking styles, C2SER first generates speech transcripts and descriptive captions of speaking styles and then infers the final speech emotion based on the aggregated context.
Explicit CoT data. We leverage signal processing tools and text LLMs to construct training data for explicit CoT. Given an emotional speech corpus that includes text transcripts and emotion labels, we extract key attributes—such as speaking rate, pitch, and energy—and map these to discrete levels (e.g., low, medium, high). Next, we employ the GLM-4-9B-Chat model Zeng et al. (2024) to generate a detailed reasoning path that combines text transcripts, extracted speech attributes, and corresponding emotion labels. This reasoning path serves as the explicit CoT training data.
@@ 3.3 Explicit Chain-of-Thought
3.4 Implicit Chain-of-Thought
Although the explicit CoT approach enables C2SER to address SER step by step using detailed intermediate representations of speech content and speaking styles, it also introduces inefficiencies during inference and increases the risk of error accumulation Yuen et al. (2024); Deng et al. (2024). To overcome these limitations, we propose a self-distillation strategy that transitions C2SER from explicit CoT to implicit CoT.

### v2

@@ Abstract
I Introduction
Figure 1: Overview of C2SER. The top path shows how standard models can “hallucinate” by generating irrelevant context. The bottom path demonstrates our two-step Chain-of-Thought (CoT) approach: first generating a detailed rationale (Explicit CoT), and then internalizing this capability for a direct and stable prediction (Implicit CoT).
@@ I Introduction
Audio is a multifaceted medium for communication, conveying speech prosody, vocal tone, and paralinguistic cues through its acoustic features. Large-scale audio-language models (ALMs) [alm-ref, AudioBench, ALM_1] have demonstrated substantial progress in understanding diverse forms of audio signals, which is crucial for advancing Artificial General Intelligence (AGI). With increasing data availability, computational power, and model size, significant strides have been made in speech signal comprehension, analysis, and reasoning, leading to more natural and human-like text responses.
Speech emotions are crucial in communication, influencing how individuals interact and respond through variations in tone, rhythm, and intensity. However, ALMs still face challenges in speech emotion recognition (SER) [alm2ser, alm2ser2], often exhibiting “hallucinations” that undermine their reliability. These hallucinations involve the generation of any information that is factually ungrounded in the provided audio signal. This issue ranges from the model fabricating plausible but unverified contextual details—such as speculating that a happy speaker is a “student celebrating university acceptance” when no such evidence exists—to more critical failures where it fundamentally misinterprets the emotion and invents a baseless justification. As illustrated in Figure 1, such a failure can lead to a cheerful utterance being labeled as “sadness” with the unsubstantiated rationale that “Maybe he is preparing for an exam.” Our work is specifically focused on steering the model’s reasoning to be strictly grounded in the acoustic and semantic evidence present in the audio, thereby mitigating these ungrounded outputs.
@@
In this work, we address the hallucination problem in ALM-based SER by incorporating detailed speech information and expanding the model’s reasoning length and depth. As illustrated in Figure 1, we introduce C2SER, a novel ALM designed to improve both the stability and accuracy of SER. C2SER integrates two critical components: contextual perception and a chain of thought (CoT), leveraging both speech content and speaking styles (e.g., speaking rate, pitch, energy) to facilitate emotion recognition. The encoder of Whisper [whisper], trained for automatic speech recognition (ASR), speech translation, and language identification, is employed for semantic perception. For acoustic perception, we introduce Emotion2Vec-S, a refined extension of Emotion2Vec [Emotion2Vec], designed to enhance the extraction of emotion-related information from audio. Emotion2Vec-S incorporates semi-supervised contrastive loss at the category level, improving emotional discrimination by combining self-supervised and semi-supervised learning.
@@
Recognizing that speech emotions are influenced by both speech content and speaking styles, such as aggressive speech characterized by a loud volume potentially indicating anger, C2SER employs a CoT [deepseekr1] training approach to incentivize reasoning capability. This approach decomposes the SER task into sequential steps: first perceiving speech content and speaking style, followed by emotion inference, with the assistance of prior context. This structured method imitates human thinking and reduces the possibility of hallucinations. To further enhance stability and prevent error propagation, especially in longer thought chains, C2SER introduces self-distillation, transferring knowledge from explicit to implicit CoT. This process helps minimize error accumulation, improving the model’s overall performance.
@@
We validate the effectiveness of C2SER through extensive experiments using multiple speech corpora and compare C2SER with state-of-the-art models, including SECap [SECap] and Qwen2-Audio [Qwen2-Audio]. To better simulate real-world context, we introduce a new SER test set, Emo-Emilia. Emo-Emilia is created through an automated labeling approach on the in-the-wild Emilia corpus and manually verified to ensure quality and diversity across various scenarios. Our experimental results demonstrate that C2SER significantly outperforms existing models on public test sets and Emo-Emilia in terms of weighted accuracy (WA), unweighted accuracy (UA), and Macro F1 score, while notably reducing hallucination-related errors. These findings highlight the potential of C2SER to provide stable and reliable emotion recognition in diverse contexts.
@@
  • We release the code, checkpoints, and test set to promote further research in the field of SER.
II Related Work
In this section, we review related work that is closely related to C2SER, dividing the discussion into two parts: audio-language models (ALMs) and speech emotion recognition (SER).
II-A Audio Language Model
@@
Large language models (LLMs) have made significant strides in natural language processing (NLP), showcasing remarkable capabilities across a variety of tasks [LLMsurvey, related_LLM, related_LLM1]. As audio is a critical medium of communication in human interactions and human-computer engagement, recent research has extended LLMs to integrate the audio modality, leading to the development of audio-language models (ALMs) [alm_AudioBench]. ALMs tackle tasks such as audio event detection, audio captioning, and speech recognition, serving as a cornerstone for comprehensive audio understanding [qwenaudio, related_kimi, related_osum].
@@
With the rapid advancements in both LLMs and the audio domain, ALMs have gained significant attention for their powerful general audio comprehension abilities [related_gama, related_flamingo2]. A typical ALM architecture consists of three core components: an audio encoder for modality-specific feature extraction, an LLM for text generation, and a projection layer to bridge the gap between the audio and text modalities. In addition to these foundational components, several studies have focused on refining ALM performance through innovative model architectures. For example, SALMONN [SALMONN] utilizes dual encoders to separately process speech and non-speech audio signals, effectively mitigating potential conflicts between different types of audio input. Other approaches have explored training strategies to enhance ALMs’ capabilities, with Qwen2-Audio [Qwen2-Audio] being a notable example. This model employs a comprehensive training pipeline that includes pretraining, supervised fine-tuning (SFT), and reinforcement learning from human feedback (RLHF).
@@
While significant progress has been made in improving the generalization and intelligence of ALMs, their performance in speech emotion recognition (SER) remains unsatisfactory, primarily due to hallucinations. SER is particularly challenging because speech emotions are inherently complex and context-dependent [alm_AIR-Bench], making it difficult for ALMs to interpret emotional states accurately. In many cases, the model may be misled by the content of the speech, leading to incorrect classifications or irrelevant responses.
II-B Speech Emotion Recognition
@@
Speech emotion is a key form of paralinguistic information that effectively conveys the speaker’s intent. Speech Emotion Recognition (SER) aims to automatically identify a speaker’s emotional state from raw audio, with applications in human–computer interaction, healthcare, and affective computing [ser_related1, ser_related2, ser_related3]. The typical SER pipeline consists of three stages: speech preprocessing, feature extraction, and emotion classification [review-mer1]. Early studies relied on manually engineered feature sets, such as MFCC, and simple neural network architectures like CNN and RNN, achieving basic performance on laboratory datasets (e.g., CREMA-D [CREMA-D], IEMOCAP [IEMOCAP]).
@@
To address the challenge of recognizing diverse emotional expressions in real-world environments, recent research has shifted towards self-supervised learning (SSL) models [ssl_ser, ssl_ser2], known for their powerful generalization capabilities. SSL models are trained on large-scale unlabeled speech data in an unsupervised manner, allowing them to extract rich, generalizable representations directly from raw speech waveforms. Popular SSL models have demonstrated significant effectiveness in extracting emotional features, serving as robust encoders for SER tasks. For example, Naini et al. [ssl_ser3] investigate four SSL models, WavLM [wavlm], wav2vec 2.0 [wav2vec2.0], HuBERT [Hubert], and Data2Vec [data2vec2.0], and evaluate their cross-domain generalization ability on different speech emotion corpora. Additionally, researchers have explored emotion-specific SSL models designed to capture emotion-relevant features. A popular approach involves fine-tuning SSL models on emotionally labeled data for specific emotional tasks. A prominent example is Emotion2Vec [Emotion2Vec], which is pre-trained on emotional data through self-supervised online distillation. Emotion2Vec uses both utterance-level and frame-level loss as supervision, demonstrating remarkable improvements in emotion recognition across different languages.
@@
Figure 2: The detailed architecture and two-stage training process of C²SER. Stage 1 (Explicit CoT): The model is trained to generate a step-by-step rationale by combining semantic features from Whisper and acoustic features from Emotion2Vec-S. Stage 2 (Implicit CoT): Through self-distillation, the model is trained to produce a direct emotional description, enhancing efficiency while preserving the reasoning capabilities.
III Method
In this section, we introduce the overall C2SER framework and then detail each of its key components. First, we provide an overview of the architecture and its two main modules. Next, we describe the contextual perception module responsible for extracting semantic and acoustic features. Finally, we present our explicit and implicit chain-of-thought reasoning schemes and explain how they are used to train and refine the model.
III-A Framework Overview
@@
More specifically, the contextual perception module comprises the following elements: a Whisper [whisper] encoder for semantic perception, Emotion2Vec-S for acoustic perception, and a connection model designed to align the feature dimensions with those required by the text LLM. Formally, given a speech waveform $X$, the Whisper encoder extract semantic representations $S={s_{1},s_{2},...,s_{N}}$ and the Emotion2Vec-S extracts acoustic representations $A={a_{1},a_{2},...,a_{M}}$ from $X$. Let $Y={y_{1},y_{2},...,y_{T}}$ be the text descriptions and $P={p_{1},p_{2},...,p_{L}}$ be the text prompts. The text LLM, parameterized by $\theta$, takes $S$ and $A$ as input and predicts $Y$ in an autoregressive manner. The overall process can be formulated as a conditional probability:
@@ III-B Contextual Perception
Emotion2Vec-S is built upon the universal speech emotion representation model, Emotion2Vec, which follows the architecture of data2vec [data2vec]. Emotion2Vec is pre-trained on open-source, unlabeled emotion data using self-supervised online distillation and has demonstrated superior performance compared to previous state-of-the-art models. It combines two main objectives to learn representations: an utterance-level loss and a frame-level loss. Specifically, the utterance-level loss uses dedicated tokens to learn a representation of the entire utterance’s emotion, while the frame-level loss focuses on predicting masked portions of the sequence, forcing the model to learn localized acoustic dependencies. However, a key limitation of Emotion2Vec is that both of these losses operate at the instance level. They lack an explicit mechanism to enforce that embeddings from different utterances of the same emotion category (e.g., two different ‘fear’ samples) should be closer to each other than to embeddings from a different but acoustically similar category (e.g., a ‘sadness’ sample). This can lead to confusion between similar emotional expressions like fear and sadness. To address this limitation, our proposed Emotion2Vec-S extends Emotion2Vec by introducing a category-level contrastive loss, which explicitly pulls embeddings of the same emotion category together while pushing apart those from different categories.
Based on the above observation, Emotion2Vec-S introduces a coarse-level supervision to Emotion2Vec. Vanilla Emotion2Vec expands Data2vec2.0 [data2vec2.0] with a fixed number of utterance tokens and is trained with $\mathcal{L}_{Utt}$ to learn the global emotion and $\mathcal{L}_{Frm}$ to learn the context emotion. Inspired by CLIP [clip], Emotion2Vec-S extends Emotion2Vec with a category-level contrastive loss $\mathcal{L}_{Cate}$. Specifically, let $G$ be the global embedding of Emotion2Vec after average pooling. Emotion2Vec-S applies a contrastive loss on $G$ by treating embeddings from utterances of the same emotion category as positive pairs and those from different categories as negative pairs. The model calculates cosine similarities between these embeddings, maximizing the similarity of positive pairs while minimizing that of negative pairs. The overall loss structure of Emotion2Vec-S is illustrated in Figure 3. The total loss function is formulated as follows:
Figure 3: The three types of losses used in Emotion2Vec-s. From top to bottom are utterance-level loss, frame-level loss, and category-level loss.
@@
III-C Explicit Chain-of-Thought
Explicit CoT reasoning enhances the ability of LLMs to handle specific tasks by detailing intermediate steps, thereby guiding the model through its reasoning process [excot1, excot2]. In C2SER, explicit CoT is employed to sequentially address the SER task. After the contextual perception module extracts detailed information regarding speech content and speaking styles, C2SER first generates speech transcripts and descriptive captions of speaking styles and then infers the final speech emotion based on the aggregated context.
Explicit CoT Data.The construction of high-quality training data for our explicit CoT framework follows a systematic, three-step process designed to ground the model’s reasoning in quantifiable acoustic evidence.
Step 1: Acoustic Attribute Extraction. We first employ signal processing tools to extract key acoustic features from each speech waveform. Specifically, we compute pitch contours using the PENN library2 and calculate the utterance-level mean to represent the speaker’s tone. Energy is extracted using the pyloudnorm library3, yielding a single value for the loudness level. The speaking rate is determined by dividing the number of phonemes in the transcript by the utterance’s duration, after trimming silences from the beginning and end of the waveform to ensure accuracy.
Step 2: Feature Discretization. To make these continuous acoustic values interpretable for the LLM, we discretize them into categorical labels. After collecting statistics for all utterances in our corpus, we calculate the mean ($\mu$) and standard deviation ($\sigma$) for each attribute. Following the principles of the Central Limit Theorem, we map the values into three levels: ‘Low’ ($\text{value}<\mu-\sigma$), ‘Medium’ ($\mu-\sigma\leq\text{value}\leq\mu+\sigma$), and ‘High’ ($\text{value}>\mu+\sigma$).
Step 3: CoT Path Generation. Finally, the discretized acoustic labels (‘Low’, ‘Medium’, ‘High’), the original speech transcript, and the ground-truth emotion label are programmatically inserted into a structured prompt. This prompt is then fed to the GLM-4-9B-Chat model, which is instructed to generate a natural-language reasoning path that first describes the speech characteristics and content, and then concludes with an emotion inference. The template for this prompt is shown in Table I.
TABLE I: Template for Explicit CoT data construction, label will be replaced with the corresponding value.
Prompt: Based on the provided speech features—including a speaking rate of <speaking rate label>, a volume level of <energy label>, and a pitch of <pitch label>— along with the text content ‘<text label>’ and the emotion emotion label, generate a natural and logical emotional description. Here is an example: ‘The speaker spoke at a <speaking rate label> pace, with a <pitch label> tone and <energy label> level: “<text label>”. Based on the analysis of speech characteristics, the emotion was inferred to be emotion label.’ Ensure including all speech features and logic of the description. Generated Example: The speaker spoke at a moderate pace, with a low-pitched tone and a soft volume: “together you sort of get this whole narrative of feedback …” Based on the speech characteristics, the emotion was inferred to be disgust, revealing a sense of resentment or aversion towards the described situation.
@@ III-C Explicit Chain-of-Thought
III-D Implicit Chain-of-Thought
Although the explicit CoT approach enables C2SER to address SER step by step using detailed intermediate representations of speech content and speaking styles, it also introduces inefficiencies during inference and increases the risk of error accumulation [imcot1, imcot2]. To overcome these limitations, we propose a self-distillation strategy that transitions C2SER from explicit CoT to implicit CoT.